# Reliable and efficient inverse analysis using physics-informed neural networks with distance functions and adaptive weight tuning

### v1

@@ 2.2 R-function-based distance function
where $\bm{\nu}$ is the inward unit normal vector on $\Gamma$ ($\bm{\nu}=-\bm{n}$). Since $\phi$ is an approximation of $\Phi$, it is desirable that it possesses similar properties. Specifically, $\phi$ is required to satisfy properties (A) and (B) as $\Phi$ does, while property (C) is relaxed to a finite order.
@@
Such a function $\phi$ is said to be normalized to the $m$-th order or an $m$-th order approximation of $\Phi$ [51, 64]. In the following, we denote $\phi=\phi^{(m)}$ to emphasize the order of normalization, if necessary.
2.2.1 Distance to a line segment
@@
Figure 1: Distance to a line segment. Representations of functions defined in Equations (9), (10), (11), and its gradient. The line segment $S_{PQ}$, realized as the intersection of $\mathbb{S}=\left\{\bm{x}\in\mathbb{R}^{2}\mid s_{PQ}\left(\bm{x}\right)=0\right\}$ and $\mathbb{T}=\left\{\bm{x}\in\mathbb{R}^{2}\mid t_{PQ}\left(\bm{x}\right)\geq 0\right\}$, is represented by =$\mathbb{P}=\left\{\bm{x}\in\mathbb{R}^{2}\mid\phi_{PQ}\left(\bm{x}\right)=0\right\}$. $\nabla\phi_{PQ}$ offers the unit normal vector in the vicinity of $S_{PQ}$.
@@
[…]
@@
where $\bm{x}_{M}=(\bm{x}_{P}+\bm{x}_{Q})/2$ is the midpoint of $S_{PQ}$. The region $\mathbb{T}=\left\{\bm{x}\in\mathbb{R}^{d}\mid t_{PQ}\left(\bm{x}\right)\geq 0\right\}$ corresponds to the circular disk whose radius is $\|\bm{x}_{PQ}\|/2$ and center is $\bm{x}_{M}$. Finally, interpreting the line segment $S_{PQ}$ as the intersection of $\mathbb{S}$ and $\mathbb{T}$, we define the sufficiently smooth approximate distance function (ADF) to the line segment $S_{PQ}$ as follows [50]:
@@
2.2.2 Distance to a domain
@@
Figure 2: Distance to a domain: square. Normalized ADFs are smooth approximations of the EDF, and approach the EDF as the normalization order is increased.
(a) $m=1$
(b) $m=2$
(c) $m=4$
(d) $m=8$
Figure 3: Distance to a domain: heart. Distance to each segment is defined as in Equation (11), then the ADF is constructed by the joining operation as in Equation (12). The ADFs are normalized to the 1st, 2nd, 4th, and 8th orders.
(a) Normalized, $\phi^{(1)}$
(b) Normalized, $\phi^{(2)}$
(c) Not normalized, $\varphi$
Figure 4: Distance to a domain: square (slice). (Top) slice along $y=0.5$. (Bottom) slice along $y=0.2$. The 1st derivative of the normalized ADFs form the inward unit normal at the boundary (property (b)), while the 1st derivative of the non-normalized ADF does not (right column). It should be noted that simply multiplying the non-normalized ADF by a constant does not help to satisfy property (b).
Assuming that the boundary $\Gamma$ is composed of piecewise segments ($\Gamma=\bigcup_{i=1}^{N}\Gamma_{i}$), the ADF to $\Gamma$ is constructed via the following joining operation [64, 57]:
@@
where $\phi_{i}$ is the ADF to $\Gamma_{i}$ ($\phi_{i}=0$ on $\Gamma_{i}$) and $m\left(\in\mathbb{Z}_{>0}\right)$ is the normalization order associated with the property (c). Equation (12) is formally known as the R-equivalence operation. While another joining, the R-conjunction operation, is also available, we employ the R-equivalence operation in this study due to its associative property, which the R-conjunction operation lacks. Figure 2 shows the ADFs with different normalization orders and the EDF to a unit square. As depicted, the ADFs serve as smooth approximations of the EDF. The ADFs $\phi^{(m)}$ approach the EDF $\Phi$ as the normalization order $m$ increases, as described in [64]. This is further confirmed by comparing $\phi^{(64)}$ and $\Phi$ in Figure 2. It should be noted that the ADF obtained by Equation (12) generalizes to curves, complex geometries (e.g., Figure 3), and higher-dimensional field [46, 64, 52, 65, 55, 56].
@@
The choice of the normalization order $m$ in Equation (12) is arbitrary, and $m=2$ is commonly used in computational mechanics practice [57]. Our preliminary numerical experiments using $m=1,2,4,8$ found that $m=1$ and $m=2$ achieve similar accuracy and convergence, while $m=4$ and $m=8$ can degrade the convergence speed (see Appendix B). Reflecting this observation and the previous studies [46, 58], we use $m=1$ for the remainder of this study.
As a simpler joining operation, one may consider the following ADF:
$\varphi\coloneq\prod_{i}\phi_{i}.$ (13)
Although Equation (13) is known to be another class of R-equivalence operations, and $\varphi$ is easy to construct and satisfies property (a) ($\varphi=0$ on $\Gamma$), it fails to meet property (b) ($\partial_{\bm{\nu}}\varphi\neq 1$ on $\Gamma$). Consequently, $\varphi$ is said to be not normalized [64]. As evidence, Figure 4 presents the ADFs and their 1st and 2nd derivatives at the slices along $y=0.5$ and $y=0.2$. It is clear that $\phi^{(1)}$ and $\phi^{(2)}$ satisfy properties (a) and (b), while $\varphi$ does not.
In this study, we employ Equation (12) to define the ADFs and the unit normal vector on the boundary instead of Equation (13), in order to address mixed boundary value problems. While distance functions in the form of $\varphi$ have been used in several studies  [39, 45, 34, 66], their applicability to complex geometries and mixed boundary value problems, where homogeneous/inhomogeneous Dirichlet, Neumann, Robin, and their combinations are prescribed, is severely limited. Specifically, methods like those in [39, 66] enforce the Dirichlet condition in a hard manner, while relying on a penalty approach for the Neumann boundary condition. This limitation arises from the difficulty in precisely defining the unit normal, which is crucial for the Neumann and mixed problems. In addition, the methods presented therein are constrained to problems defined on convex hulls, and cannot be extended to form distance functions to non-convex geometries, such as shown in Figure 3 and discussed in Section 3.
2.2.3 Approximate solution structure
@@
For simplicity, we only discuss the treatment of Dirichlet boundary condition. However, the method can be extended to Neumann, Robin, and mixed boundary conditions, as described in [46, 55, 56, 67] and demonstrated in Section 3.
As mentioned in Section 1, boundary conditions can be imposed in either soft [9] or hard manner [2, 39, 46], in the context of PINN. In the soft imposition approach, the raw output of the neural network serves as the approximate solution:
@@
where $\bar{g}_{D}$ is an interpolation of the Dirichlet condition ($\bar{g}_{D}=g_{D}$ on $\Gamma_{D}$), and $\phi_{D}$ is the approximate distance function (ADF) to $\Gamma_{D}$. With $\phi_{D}$ vanishing on $\Gamma_{D}$, $\tilde{u}$ readily satisfies the imposed Dirichlet boundary condition. When the Dirichlet boundary $\Gamma_{D}$ consists of multiple segments $\Gamma_{i}$ ($i=1,\ldots,N_{D}$) (i.e., Dirichlet conditions are prescribed on multiple segments), the interpolant $\bar{g}_{D}$ can be constructed by transfinite interpolation [51]:
@@
2.3 Adaptive weight tuning for gradient imbalance
The formulation described in Section 2.2 ensures that the approximate solution inherently satisfies the prescribed boundary conditions. Consequently, $\mathcal{L}_{\text{BC}}$ in Equation (5) becomes identically zero. This implies that, for forward problems, minimizing $\mathcal{L}_{\text{PDE}}$ alone is sufficient, akin to classical numerical methods. However, in inverse problems, PINN must incorporate observed data through Equation (8). In this case, the loss function with hard boundary condition enforcement takes the form:
@@ 2.3 Adaptive weight tuning for gradient imbalance
where ~$\tilde{\mathcal{L}}_{\text{PDE}}$ and ~$\tilde{\mathcal{L}}_{\text{Data}}$ are the modified PDE and data loss functions, respectively, obtained by replacing $\hat{u}$ with $\tilde{u}$ in Equations (6) and (8).
A critical challenge in optimizing multi-objective functions, such as Equation (22), is the risk of convergence to undesired, suboptimal Pareto solutions due to gradient imbalance, which can further lead to the emergence of trivial, constant solutions, known as failure modes [68, 63, 47, 69]. In such cases, appropriate normalization is crucial to balance the gradients of multiple loss terms. In the context of PINN, the PDE and data loss terms are typically of different magnitudes, and their gradients may not be directly comparable. To address this issue, we employ an adaptive weight tuning method known as dynamic normalization [47]. This method adaptively adjusts the weight $\lambda_{\text{Data}}$ in Equation (22) to balance the gradients of each loss term during the process of optimization, eliminating the need for manual tuning.
@@
Here, $\Delta\tilde{\mathcal{L}}_{\text{PDE}}$ and $\Delta\tilde{\mathcal{L}}_{\text{Data}}$ denote the changes in the PDE and data loss terms, respectively. For simplicity, we assume that each gradient is orthogonal to the other. While this assumption may not always hold, it greatly simplifies the scope, and the resultant adaptive weights have been shown to provide significant performance improvements in practice  [30, 24, 70, 47, 71]. If the orthogonality assumption poses a severe issue, projection techniques such as those presented in [72] can be employed to mitigate conflicts by projecting backpropagated gradients onto orthogonal directions. Under the aforementioned orthogonality assumption, we split Equation (25) into:
@@
Based on the analysis in [30, 63], it is desirable to balance the reduction of the PDE and data loss such that they are reduced at similar rates, $\Delta\tilde{\mathcal{L}}_{\text{PDE}}\sim\Delta\tilde{\mathcal{L}}_{\text{Data}}$. This translates to the following adaptive weight tuning strategy:
@@
In practice, the exponential decay is applied to stabilize the fluctuations  [24, 30, 31]:
@@
where $\beta\in[0,1)$ is the decay rate. Our preliminary numerical experiments indicate that $\beta$ should be set to a value close to 1 to ensure stability. However, as is well known, setting $\beta$ too close to 1 can lead to slow convergence. As an illustration, applying Equation (29) repeatedly, one obtains:
@@
One can instantly realize that $\hat{\lambda}_{\text{Data}}^{(n)}$ is biased towards its initial value $\hat{\lambda}_{\text{Data}}^{(0)}$, particularly when $\beta$ is close to 1 and $n$ is not sufficiently large. To remove this initialization bias, we set $\hat{\lambda}_{\text{Data}}^{(0)}=0$ and introduce bias correction as in [73]:
@@
ergo $\mathbb{E}[\tilde{\lambda}_{\text{Data}}^{(n)}]=\mathbb{E}[\lambda_{\text{Data}}^{(n)}]$, whereas $\mathbb{E}[\hat{\lambda}_{\text{Data}}^{(n)}]\neq\mathbb{E}[\lambda_{\text{Data}}^{(n)}]$. Equation (32) ensures that the weight is properly adjusted, even in the early stages of optimization and when $\beta$ is near 1. Despite the presence of numerous studies on adaptive weight tuning in the context of PINN  [30, 24, 70, 31], discussion on this initialization bias is often overlooked. Bias correction technique in Equation (32) is crucial for the appropriate adjustment, ensuring $\Delta\tilde{\mathcal{L}}_{\text{PDE}}\sim\Delta\tilde{\mathcal{L}}_{\text{Data}}$ is satisfied from the initial stage. The effectiveness of bias-corrected dynamic normalization has already been demonstrated in [47, 74] for both forward and inverse problems, and is further confirmed in Section 3.
2.4 Positivity enforcement for inverse analysis
When applied to inverse problems, PINNs are typically equipped with additional trainable parameters to represent the physical quantities of interest. For instance, when identifying the diffusion coefficient $\kappa$ in a diffusion equation, an additional parameter $\hat{\kappa}$ is introduced to represent and estimate $\kappa$ from the available observations. This approach has been widely used in the literature, as exemplified in [9, 14, 75, 76]. Since $\hat{\kappa}$ is a trainable parameter, it is updated alongside the network parameters (weights and biases). This allows $\hat{\kappa}$ to take any real value, including negative values. Such an approach can be viewed as a class of soft imposition, as the sign of $\hat{\kappa}$ is not explicitly constrained and is only expected to be learned through the optimization process. From a physical point of view, however, these quantities must possess specific signs to ensure physical validity. In the case of diffusion example, the diffusion coefficient $\kappa$ must remain positive.

### v3

@@ 2.2.1 R-function-based distance fields
where $\bm{\nu}$ is the inward unit normal vector on $\Gamma$ ($\bm{\nu}=-\bm{n}$). Since $\phi$ is an approximation of $\Phi$, it is desirable that it possesses similar properties. Specifically, $\phi$ is required to satisfy properties (A) and (B) as $\Phi$ does, while property (C) is relaxed to a finite order:
@@
Such a function $\phi$ is said to be normalized to the $m$-th order or an $m$-th order approximation of $\Phi$ [Rvachev2001, Biswas2004]. In the following, we denote $\phi=\phi^{(m)}$ to emphasize the order of normalization, if necessary.
Distance to a line segment
@@
Figure 1: Distance to a line segment. Representations of functions defined in Equations (9), (10), (11), and its gradient. The line segment $S_{PQ}$, realized as the intersection of $\mathbb{S}=\left\{\bm{x}\in\mathbb{R}^{2}\mid s_{PQ}\left(\bm{x}\right)=0\right\}$ and $\mathbb{T}=\left\{\bm{x}\in\mathbb{R}^{2}\mid t_{PQ}\left(\bm{x}\right)\geq 0\right\}$, is represented by $\mathbb{P}=\left\{\bm{x}\in\mathbb{R}^{2}\mid\phi_{PQ}\left(\bm{x}\right)=0\right\}$. $\nabla\phi_{PQ}$ offers the unit normal vector in the vicinity of $S_{PQ}$.
@@
where $\|\cdot\|$ denotes the $L^{2}$-norm, $s^{\prime}_{PQ}$ is the signed distance to the infinite line $S^{\prime}_{PQ}$, $\bm{x}_{PQ}=\bm{x}_{Q}-\bm{x}_{P}$ is the vector from $P$ to $Q$, and $\bm{n}_{PQ}$ is the unit normal to $S^{\prime}_{PQ}$. The function $s_{PQ}$ can be viewed as the signed distance function to the infinite line $S^{\prime}_{PQ}$, normalized by the length of the segment of interest, $S_{PQ}$. With these definitions, the line segment $S_{PQ}$ is realized as a subset of $\mathbb{S}=\left\{\bm{x}\in\mathbb{R}^{d}\mid s_{PQ}\left(\bm{x}\right)=0\right\}$. Next, we introduce the trimming function (TF) $t_{PQ}$:
@@
where $\bm{x}_{M}=(\bm{x}_{P}+\bm{x}_{Q})/2$ is the midpoint of $S_{PQ}$. The region $\mathbb{T}=\left\{\bm{x}\in\mathbb{R}^{d}\mid t_{PQ}\left(\bm{x}\right)\geq 0\right\}$ corresponds to the circular disk whose radius is $\|\bm{x}_{PQ}\|/2$ and center is $\bm{x}_{M}$. Finally, interpreting the line segment $S_{PQ}$ as the intersection of $\mathbb{S}$ and $\mathbb{T}$, we define the sufficiently smooth approximate distance function (ADF) to the segment $S_{PQ}$ as follows [Sheiko1982]:
@@
Distance to the boundary: normalized approach
@@
Figure 2: Distance to the boundary: square. Normalized ADFs are smooth approximations of the EDF, and approach the EDF as the normalization order is increased.
(a) Naïve, $\psi$
(b) Non-normalized, $\varphi$
(c) Normalized, $\phi^{(1)}$
(d) Normalized, $\phi^{(8)}$
Figure 3: Distance to the boundary: non-convex shapes. (Top) L-shaped geometry. (Bottom) heart-shaped geometry. Colors represent the distance fields, and arrows indicate the gradients (arrow lengths are scaled only in a relative sense in each panel). The naïve approach (first column) fails to capture the non-convex geometry with undesired zero level sets emerging inside the domain. The non-normalized formulation (second column) is able to represent non-convex domains, but the gradients are not normalized (i.e., they do not have unit length on the boundary). The normalized formulations (third and fourth columns) accurately capture the geometry while providing unit normals at the boundary.
(a) Naïve, $\psi$
(b) Non-normalized, $\varphi$
(c) Normalized, $\phi^{(1)}$
(d) Normalized, $\phi^{(2)}$
Figure 4: Distance to the boundary: square (slice). (Top) slice along $y=0.5$. (Bottom) slice along $y=0.2$. The first derivative of the normalized ADFs form the inward unit normal at the boundary (property (b)), whereas the naïve and non-normalized ADFs fail to do so. Simply scaling the naïve or non-normalized ADF ($\psi$ or $\varphi$) by a constant does not resolve this issue.
Assuming that the boundary $\Gamma$ is composed of piecewise segments ($\Gamma=\bigcup_{i}\Gamma_{i}$), the normalized ADF to $\Gamma$ is constructed via the following joining operation [Biswas2004, Millan2015]:
@@
where $\phi_{i}$ is the ADF to $\Gamma_{i}$ ($\phi_{i}=0$ on $\Gamma_{i}$, as defined in Equation (11)), and $m\left(\in\mathbb{Z}_{>0}\right)$ is the normalization order associated with the property (c). Equation (12) is formally known as the R-equivalence operation. While another joining, the R-conjunction operation, is also available, we employ the R-equivalence operation in this study due to its associative property, which the R-conjunction lacks. Figure 2 shows the ADFs with different normalization orders and the EDF to a unit square. As depicted, the ADFs serve as smooth approximations of the EDF. The ADFs $\phi^{(m)}$ approach the EDF $\Phi$ as the normalization order $m$ increases, as described in [Biswas2004]. This is further confirmed by comparing $\phi^{(64)}$ and $\Phi$ in Figure 2. It is also noteworthy that the ADF obtained by Equation (12) generalizes to curves [Biswas2004], complex non-convex geometries [Shapiro2002SAGE, Tsukanov2003MeshfreeFluid], and higher-dimensional field such as 3D objects [Biswas2004, Shapiro2007] and 4D cubes [Sukumar2022MixedBC].
@@
The choice of the normalization order $m$ in Equation (12) is arbitrary, and $m=2$ is commonly used in computational mechanics practice [Millan2015]. Our preliminary numerical experiments using $m=1,2,4,8$ found that $m=1$ and $m=2$ achieve similar accuracy and convergence, while $m=4$ and $m=8$ can degrade the convergence speed (see Appendix B). Reflecting this observation and the previous studies [Sukumar2022MixedBC, Berrone2023BCEnforcement], we use $m=1$ for the remainder of this study.
Distance to the boundary: potential alternatives
As outlined in Section 1, several alternatives could be considered to realize the distance to the boundary $\Gamma=\bigcup_{i}\Gamma_{i}$. In particular, the community has favored the naïve approach ((3-i), as introduced in Section 1), under the assumption that the domain is convex (see, e.g., [Lagaris1998, Lu2021HardDesign, Li2024HardAdvcDiff]). In this formulation, the domain is detected as the intersection of half-spaces, each of which represented by the SDF to an infinite line (or plane) containing $\Gamma_{i}$. For instance, in the case of a unit square, the domain is represented by the intersection of four half-spaces:
$\psi\coloneq\prod_{i}s_{i},$ (13)
where $\psi$ is the naïve ADF, and each $s_{i}$ is defined as in Equation (9), namely $s_{1}=x$, $s_{2}=1-x$, $s_{3}=y$, and $s_{4}=1-y$. This formulation can be extended to other polygons and polyhedra, as long as they are convex. Nonetheless, it is insufficient for non-convex geometries, as it does not account for segment endpoints (or plane edges). Figure 3 (a, top) illustrates the distance field generated by Equation (13) for an L-shaped domain, where $s_{5}=0.5-x$ and $s_{6}=0.5-y$ are included to the previous four $s_{i}$. One can observe that false boundary detections (zero level sets) appear inside the domain, due to the presence of a concave corner (specifically, $s_{5}$ and $s_{6}$). The corresponding gradients also exhibit inconsistent scale and direction on the boundary, violating property (b). We further applied this approach to represent a heart-shaped geometry (Figure 3 (a, bottom)), by defining SDFs for the top humps and the bottom cusp separately, and combining them through Equation (13). Despite this careful construction, the resulting distance field shows a sudden jump inside the domain, undefined regions outside the arcs, and undesired boundary detections. In addition to these geometric shortcomings, applying this naïve formulation to Neumann and mixed conditions remain challenging due to the lack of a well-defined unit normal on the boundary. These observations suggest that while the naïve approach (3-i) may suffice for convex domains with relatively simple BCs, such as homogeneous Dirichlet conditions, it is unsuitable for complex geometries and a wide range of boundary value problems.
As another alternative, one may consider:
$\varphi\coloneq\prod_{i}\phi_{i},$ (14)
where $\varphi$ denotes the non-normalized ADF, and each $\phi_{i}$ is defined in Equation (11). By using $\phi_{i}$ instead of $s_{i}$, this formulation is aware of segment endpoints (plane edges) and can therefore represent non-convex geometries, as shown in Figure 3 (b). Equation (14) is known as another class of R-equivalence operation, and $\varphi$ satisfies $\varphi=0$ on $\Gamma$ (property (a)). However, it should be noted that this ADF does not satisfy property (b) ($\partial_{\bm{\nu}}\varphi\neq 1$ on $\Gamma$, see Figure 3 (b, top)), and is therefore called non-normalized in the framework of R-functions [Biswas2004]. In summary, this non-normalized formulation (3-ii) offers greater geometric flexibility than the naïve approach (3-i), but remains limited in enforcing various boundary conditions.
To further assess whether the ADFs in Equations (13), (14), and (12) satisfy properties (a), (b), and (c), Figure 4 compares naïve, non-normalized, and normalized distance fields (to the first and second orders), and their derivatives for a unit square (slices along $y=0.5$ and $y=0.2$ are shown). Both $\phi^{(1)}$ and $\phi^{(2)}$ (formulation (3-iii), Equation (12)) satisfy properties (a) and (b), with $\phi^{(2)}$ also satisfying (c) up to the second order. In contrast, $\psi$ ((3-i), Equation (13)) and $\varphi$ ((3-ii), Equation (14)) only meet property (a). Moreover, from Figure 4 (a) and (b), one can see that simply scaling $\psi$ or $\varphi$ by a constant, as done in [Lu2021HardDesign], does not help to enforce property (b).
Based on these observations, we employ Equation (12) in this study to define the ADFs and unit normal vectors on the boundary, rather than Equations (13) or (14), in order to address mixed boundary value problems. Therefore, we use the term ‘hard imposition’ to refer specifically to the normalized formulation, i.e., Equation (12) (approach (3-iii)) in the remainder of this paper. This choice becomes further beneficial for imposing inhomogeneous Dirichlet and mixed boundary conditions, as discussed in the following section.
Approximate solution structure
@@
For simplicity, we only discuss the treatment of Dirichlet boundary condition. However, the method can be extended to Neumann, Robin, and mixed boundary conditions, as described in [Sukumar2022MixedBC, Shapiro2002SAGE, Tsukanov2003MeshfreeFluid, Rvachev2000] and demonstrated in Section 3.
As mentioned in Section 1, boundary conditions can be imposed in either soft [Raissi2019PINN] or hard manner [Lagaris1998, Sun2020Surrogate, Sukumar2022MixedBC], in the context of PINN. In the soft imposition approach, the raw output of the neural network serves as the approximate solution:
@@
where $\bar{g}_{D}$ is an interpolation of the Dirichlet condition ($\bar{g}_{D}=g_{D}$ on $\Gamma_{D}$), and $\phi_{D}$ is the approximate distance function (ADF) to $\Gamma_{D}$. With $\phi_{D}$ vanishing on $\Gamma_{D}$, $\tilde{u}$ readily satisfies the imposed Dirichlet boundary condition. When the Dirichlet boundary $\Gamma_{D}$ consists of multiple segments $\Gamma_{i}$ ($i=1,\ldots,N_{D}$) (i.e., Dirichlet conditions are prescribed on multiple segments), the interpolant $\bar{g}_{D}$ can be constructed by transfinite interpolation [Rvachev2001]:
@@
2.2.2 Adaptive weight tuning for gradient imbalance
The formulation described in Section 2.2.1 ensures that the approximate solution inherently satisfies the prescribed boundary conditions. Consequently, $\mathcal{L}_{\text{BC}}$ in Equation (5) becomes identically zero. This implies that, for forward problems, minimizing $\mathcal{L}_{\text{PDE}}$ alone is sufficient, akin to classical numerical methods. However, in inverse problems, PINN must incorporate observed data through Equation (8). In this case, the loss function with hard boundary condition enforcement takes the form:
@@ 2.2.2 Adaptive weight tuning for gradient imbalance
where $\tilde{\mathcal{L}}_{\text{PDE}}$ and $\tilde{\mathcal{L}}_{\text{Data}}$ are the modified PDE and data loss functions, respectively, obtained by replacing $\hat{u}$ with $\tilde{u}$ in Equations (6) and (8).
A critical challenge in optimizing multi-objective functions, such as Equation (23), is the risk of convergence to undesired, suboptimal Pareto solutions due to gradient imbalance, which can further lead to the emergence of trivial, constant solutions, known as failure modes [Krishnapriyan2021FailureModes, Rohrhofer2023IEEE, Deguchi2023DynNorm, Daw2023PropagationFailures]. In such cases, appropriate normalization is crucial to balance the gradients of multiple loss terms. In the context of PINN, the PDE and data loss terms are typically of different magnitudes, and their gradients may not be directly comparable. To address this issue, we employ an adaptive weight tuning method known as dynamic normalization [Deguchi2023DynNorm]. This method adaptively adjusts the weight $\lambda_{\text{Data}}$ in Equation (23) to balance the gradients of each loss term during the process of optimization, eliminating the need for manual tuning.
@@
Here, $\Delta\tilde{\mathcal{L}}_{\text{PDE}}$ and $\Delta\tilde{\mathcal{L}}_{\text{Data}}$ denote the changes in the PDE and data loss terms, respectively. For simplicity, we assume that each gradient is orthogonal to the other. This assumption may not strictly hold, as $\bm{\theta}$ is shared across multiple objectives and gradient correlations can arise. Nonetheless, it offers a tractable simplification to facilitate analysis, and prior studies have shown that correcting magnitude imbalances typically plays a more critical role in practice [Wang2021GradPath, Maddu2022InvDirichlet, Rohrhofer2021arXiv]. If directional conflicts become pronounced or pose significant challenges in optimization, projection techniques such as those presented in [Yu2020PCGrad, Li2023AGDA] can be employed to mitigate them. Under the aforementioned orthogonality assumption, we split Equation (26) into:
@@
Based on the analysis in [Wang2021GradPath, Rohrhofer2023IEEE], it is desirable to balance the reduction of the PDE and data loss such that they are reduced at similar rates, i.e., $\Delta\tilde{\mathcal{L}}_{\text{PDE}}\sim\Delta\tilde{\mathcal{L}}_{\text{Data}}$. This translates to the following adaptive weight tuning strategy:
@@
In practice, the exponential decay is applied to stabilize the fluctuations [Maddu2022InvDirichlet, Wang2021GradPath, Wang2023ExpertGuide]:
@@
where $\beta\in[0,1)$ is the decay rate. Our preliminary numerical experiments indicate that $\beta$ should be set to a value close to 1 to ensure stability. However, as is well known, setting $\beta$ too close to 1 can lead to slow convergence. As an illustration, applying Equation (30) recursively, one obtains:
@@
One can instantly realize that $\hat{\lambda}_{\text{Data}}^{(n)}$ is biased towards its initial value $\hat{\lambda}_{\text{Data}}^{(0)}$, particularly when $\beta$ is close to 1 and $n$ is not sufficiently large. To remove this initialization bias, we set $\hat{\lambda}_{\text{Data}}^{(0)}=0$ and introduce bias correction as in [Kingma2015Adam]:
@@
ergo $\mathbb{E}[\tilde{\lambda}_{\text{Data}}^{(n)}]=\mathbb{E}[\lambda_{\text{Data}}^{(n)}]$, whereas $\mathbb{E}[\hat{\lambda}_{\text{Data}}^{(n)}]\neq\mathbb{E}[\lambda_{\text{Data}}^{(n)}]$. Equation (33) ensures that the weight is properly adjusted, even in the early stages of optimization and when $\beta$ is near 1. Despite the presence of numerous studies on adaptive weight tuning in the context of PINN  [Wang2021GradPath, Maddu2022InvDirichlet, Jin2021NSFnets, Wang2023ExpertGuide], discussion on this initialization bias is often overlooked. Bias correction technique in Equation (33) is crucial for the appropriate adjustment, ensuring that $\Delta\tilde{\mathcal{L}}_{\text{PDE}}\sim\Delta\tilde{\mathcal{L}}_{\text{Data}}$ is satisfied from the initial stage. The effectiveness of bias-corrected dynamic normalization has already been demonstrated in [Deguchi2023DynNorm, Deguchi2023Keisan] for both forward and inverse problems, and is further confirmed in Section 3.
2.2.3 Positivity enforcement for inverse analysis
When applied to inverse problems, PINNs are typically equipped with additional trainable parameters to represent the physical quantities of interest. For instance, when identifying the diffusion coefficient $\kappa$ in a diffusion equation, an additional parameter $\hat{\kappa}$ is introduced to represent and estimate $\kappa$ from the available observations. This approach has been widely used in the literature, as exemplified in [Raissi2019PINN, Kharazmi2021hpVPINN, Jagtap2020CPINN, Lu2021DeepXDE]. Since $\hat{\kappa}$ is a trainable parameter, it is updated alongside the network parameters (weights and biases). This allows $\hat{\kappa}$ to take any real value, including negative values. Such an approach can be viewed as a class of soft imposition, as the sign of $\hat{\kappa}$ is not explicitly constrained and is only expected to be learned through the optimization process. From a physical point of view, however, these quantities must possess specific signs to ensure physical validity. In the case of diffusion example, the diffusion coefficient $\kappa$ must remain positive.